Problem 2 ENVL GEOL 3435 | Spring 2019 At a water elevation of 6391 ft, Mono Lake has a volume of 2,939,000 ac-ft, and a surface area of 48,10o ac. Annual inputs to the lake include 8 in. of direct precipitation, runoff fróm gauged streams of 150,000 ac-ft per year, and ungauged runoff and groundwater inflow of 37,000 ac-ft per year. Evaporation is 49 in. per year. Atotal of 85,000 ac-ft of wate rise or fall? Provide evidence for your answer r are diverted from the basin each year. Will the lake level Answer: Work: ring 2019 çno La has a volume of 2,939,000
Answer #1

Total volume of lake Mono= 2,939,000 ac-ft

There are two types of feedback to the lake : 1) Positive feedback and 2) Negative feedback

Positive ,adding water to the lake =

Rainfall =8 inch per year = 0.667 ft year

Volume added to lake = Surface area of lake* rainfall = 48,100 *0.667 = 32,082.7 ac-ft per year

Runoff from stream = 150,000 ac-ft per year

Ground water baseflow = 37,000 ac-ft per year

Adding these three = (32,082.7+150,000+37,000) ac ft per year = 219.082.7 ..................... (1)

Water going out from lake

Evaporation = 45 inch .Per year = 3.75 feet per year

Volume lost by evaporation = surface area of lake * evaporation = 48,100* 3.75=180, 375 ac-ft per year

Water diverted to other basin = 85,000 ac feet per year

Total water lost from lake = (180,3 75 + 85,000)ac-ft per year

=265,375 ac-ft per year

Net volume of water in lake as per water balance eq

= Volume of lake ( water added to lake - water lost from lake )

=2,939,000+(219,082.7-265,375) ac-ft per year

= 2,939,000(-46,292.3)= 2,939,0 00-46,292.3

=2,892,707.7 ac-ft

Hence as there is found to be a total decrease in the volume of lake ,hence water level in Mono lake will fall at the end .